40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2074     Entry time: Fri Oct 9 03:53:56 2009
Author: Jenne 
Type: Update 
Category: Adaptive Filtering 
Subject: Remaking the ASS 

The c1ass computer, which is now used for the OAF system, has many remnants from the days when it was actually used as an ASS.  These PIT and YAW filter banks and other things were taking up a lot of unnecessary space, so I deleted them in the ass.mdl file.  These files are all backed up, so we can always revert back to an older version when we want some Alignment Stabilization again someday.  I then did a make ass, following the instructions on the 40m Wiki -> Computers and Scripts -> Simulink to Front-End Code page.  Rana moved some things around, most notably all of the things (like the ASS screens) which were only in ...../users/alex/.... are now in ....../caltech/cds/advLigo/..... .  This required a few restarts of the c1ass machine (after a couple different versions of the simulink diagram....one to make sure we knew how to do it, and then again actually deleting the unused portions).

The big lesson of the night was that there are 2 signal paths for the PEM channels.  As is shown in Figure 3 in the mevans document, the PEM channels get the matching filters when they go to the adaptation algorithm, but when they go to the FIR filter, they do not get the matching filters. This is implemented by taking the output of the giant PEM matrix, and having a duplicate of each of the channels "selected for adaptation", one which gets filtered through the PEM_N_ADPT banks, and one which goes straight (in code-land) to the FIR filter.  So, it seems like all the filters which we had been including in the input side of the matrix for matching purposes need to be put in the output side.  One of the AA32 filters needs to stay in the input side, for actual anti imaging of the PEM channels, then we put the AA32 and AI32 which are for matching the ERR_EMPH and CORR filter banks up in the PEM_N_ADAPT banks.  Rana and I made these filters, and they are now turned on appropriately with the OAF down script (so that all the filters are ready and waiting for the OAF to be turned on).

A little success with getting the 3Hz peak reduced, but not a lot beyond that.  Tomorrow I'll put the accelerometers back where they used to be to see if they help out at all.

ELOG V3.1.3-